Hello i'm stuck here and have been trying to get it but no luck it's
21 5/9 -7 11/15

Answer by JulietG(1812)

Solve the same way as the other.
Find the lowest common denominator of the fraction
9,18,27,36,45,54...
15,30,45,60...
Use 45
21 5/9 = 21 25/45
7 11/15 = 7 33/45
21 wholes + 25/45 = 21*45 945+25 = 970/45
7 wholes + 33/45 = 7*45 315+33 = 348/45
subtract
970/45 - 348/45 = 655/45
Reduce
14 25/55
Reduce the fraction
14 5/11